# Towards Understanding Variants of Invariant Risk Minimization through the Lens of Calibration

### v3

@@ 6 Information Bottleneck in Calibration
In our experiments, we observe that IRM with information bottleneck is effective in calibration, as shown in Figure 2 and Figure 5(c). Therefore, in this section, we provide the results of the ablation study for IBIRM.
@@
First, we investigated the impact of the penalty term related to the information bottleneck in IBIRM on calibration. Figure 6 visualizes various evaluation metrics in CMNIST when altering the strength of the information bottleneck in IBIRM, that is, the magnitude of $\gamma$ in eq. 7. As shown in panel (a), which visualizes this based on OOD ECE and OOD Accuracy, we confirm the clear trend that strengthening the regularization of $\gamma$ improves both accuracy and ECE. Panel (b) shows the relationship between ECE in ID and OOD. Increasing $\gamma$ aligns both ECEs towards the red line, representing equality for conditional probability. In other words, this constitutes a successful calibration across environments. These findings suggest that not only is the IRMv1’s regularization term aimed at OOD scenarios, but the information bottleneck also further promotes the learning of invariant features.
@@
Additionally, we demonstrate the validity of our results based on the theoretical principles of the information bottleneck. In OOD scenarios, the objective of the information bottleneck is to minimize the mutual information $I(X;\Phi(X))$ between the input $X$ and the intermediate representation $\Phi(X)$ while preserving the mutual information $I(\Phi(X);Y)$ between $\Phi(X)$ and the output $Y$. Intuitively, this means learning $\Phi(X)$ to reduce as much information from $X$ as possible while retaining as much information about $Y$. When $\Phi(X)$ is a deterministic transformation of $X$, the entropy $H(\Phi(X))$ can be used to minimize $I(X;\Phi(X))$ (Kirsch et al., 2021). In the context of IRM, which aims to learn invariant features, keeping $H(\Phi(X))$ small encourages $\Phi$ to capture invariant features $X_{inv}$ from $X$ and discard spurious features $X_{sup}$. It should be noted that Ahuja et al. (2022) theoretically demonstrated, using realistic SEMs, that minimizing $H(\Phi(X))$ in IRM leads to $\Phi$ discarding spurious features and grasping invariant features. Our results empirically support the effectiveness of applying the information bottleneck method to invariant feature learning.
@@
[linecolor=white, outerlinewidth=1pt, roundcorner=5pt] ID-ECE = OOD-ECE
Figure 6: The figures illustrate the impact of the information bottleneck on calibration with IBIRM on CMNIST task. Panel (a) displays the relationship between ECE and Accuracy in the OOD context, investigated by varying the coefficient $\gamma$ of the information bottleneck penalty in the IBIRM formulation (eq. 7). As the value of $\gamma$ increases, both metrics show improvement. Panel (b) visualizes the values of ECE in both ID and OOD. Same as (a), altering the value of $\gamma$ shows that the larger the value, the more the data aligns with the red line, which indicates equality between the two ECEs. With a sufficiently large $\gamma$, the points are almost perfectly distributed along the red line, indicating successful calibration across multiple environments.
@@ 7 Discussion and Conclusion
In this paper, we focus on IRM, which is challenging to implement due to the difficulties of bi-level optimization. We conducted comparative evaluations using ECE, a computationally feasible metric for calibration, to understand its approximation methods better and measure how much the model learned environmental invariant features.
Our experimental results suggested that IB-IRM, which applies the information bottleneck method to IRM, is superior in calibration when we compared models with almost the same accuracy (Figure 2 and Table 2).
Another finding from our empirical results is that regularization penalties introduced for OOD generalization effectively achieve lower ECE. However, in practical datasets (under diversity shift), we observe a trade-off between OOD ECE and accuracy. Although IB-IRM mitigates this tradeoff compared to IRMv1 and other variants; carefully balancing these two metrics is important in aiming for OOD generalization.
@@ A.1.1 IRMv1
The learning process was conducted using the loss calculated based on the formulation shown in 6. The hyperparameter $\lambda$ is used to adjust the impact of the IRMv1 penalty.
@@ A.1.2 IBIRM
The learning process was performed using the loss calculated based on the formulation shown in 7. The hyperparameters $\lambda$ and $\gamma$ are used to adjust the influence of the IRMv1 penalty and the information bottleneck penalty, respectively.
@@ Appendix B Additional results
B.1 Impact of penalties of IBIRM
Figure 3 in CMNIST compares the OOD ECE of the model when varying the regularization penalty $\lambda$ in IRMv1. Additionally, it compares the OOD ECE of the model when varying the coefficients of the two regularization penalties in IBIRM 7. Similar to 3, as the regularization penalty increases, the OOD ECE decreases.
@@ B.1 Impact of penalties of IBIRM
Figure 7: Comparison of the model’s OOD ECE in CMNIST, with various values of $\gamma$ and $\lambda$ fixed at a sufficiently large value (10000) in the formulation of IBIRM 7. It was observed that as $\gamma$ increases, i.e., as the regularization penalty from the information bottleneck method becomes more significant, the ECE of the model in OOD scenarios decreases.
@@
Figure 8: Comparison of the model’s OOD ECE in CMNIST, with various values of $\lambda$ and $\gamma$ fixed at a sufficiently large value (10000) in the formulation of IBIRM 7. It was observed that as $\lambda$ increases, i.e., as the regularization penalty from IRMv1 in IB-IRM becomes more significant, the ECE of the model in OOD decreases.
B.2 Relationship between ID and OOD accuracy in diversity shift
@@ Appendix C Additional discussion and potential future directions
The experimental results demonstrated that IBIRM, which employs an information bottleneck, excels in terms of model calibration. This can be attributed to the fact that applying the information bottleneck restricts the representational power of the featurizer $\Phi(X)$, limiting its complexity and enabling it to discard spurious features, thereby reducing dependence on them. This approach of constraining the complexity of data representations has been applied in various contexts, and there are related studies in the field of Large Language Models (LLMs) as well.

### v4

@@ 6 Information Bottleneck in Calibration
In our experiments, we observe that IRM with information bottleneck is effective in calibration, as shown in Figure 2(c) and Figure 4(c). Therefore, in this section, we provide the results of the ablation study for IB-IRM.
@@
We investigated the impact of the penalty term related to the information bottleneck in IB-IRM on calibration. Figure 5(f) visualizes various evaluation metrics in CMNIST when altering the strength of the information bottleneck in IB-IRM, that is, the magnitude of $\gamma$ in eq. 9. As shown in the top row of Figure 5(f), which visualizes this based on OOD calibration and OOD accuracy, we confirm the clear trend that strengthening the regularization of $\gamma$ improves both accuracy and ECE. The bottom row of Figure 5(f) shows the relationship between the calibration metrics in ID and OOD. Increasing $\gamma$ aligns both calibration metrics towards the red line, representing that this constitutes a successful calibration across environments. These findings suggest that not only is the IRMv1’s regularization term aimed at OOD scenarios, but the information bottleneck also further promotes the learning of invariant features.
As an intuitive explanation of the information bottleneck approach’s benefit for ECE is as follows:
  1. 1.
    The information bottleneck explicitly regularizes entropy.
  2. 2.
    This leads to the prevention of overfitting to one-hot encoded labels.
  3. 3.
    As a result, the model trained by information bottleneck is well calibrated
It is known that the reason for the degradation of calibration performance in recent large-scale neural networks is due to the overconfidence caused by their complexity (Guo et al., 2017). By keeping entropy low and maintaining the model relatively simple, overconfidence is prevented, which explains the superior calibration performance of the information bottleneck.
@@
As a technical explanation, we describe the validity of our results based on the theoretical principles of the information bottleneck. In OOD scenarios, the objective of the information bottleneck is to minimize the mutual information $I(X;\Phi(X))$ between the input $X$ and the intermediate representation $\Phi(X)$ while preserving the mutual information $I(\Phi(X);Y)$ between $\Phi(X)$ and the output $Y$. Intuitively, this means learning $\Phi(X)$ to reduce as much information from $X$ as possible while retaining as much information about $Y$. When $\Phi(X)$ is a deterministic transformation of $X$, the entropy $H(\Phi(X))$ can be used to minimize $I(X;\Phi(X))$ (Kirsch et al., 2021). In the context of IRM, which aims to learn invariant features, keeping $H(\Phi(X))$ small encourages $\Phi$ to capture invariant features $X_{inv}$ from $X$ and discard spurious features $X_{sup}$. It should be noted that Ahuja et al. (2022) theoretically demonstrated, using realistic SEMs, that minimizing $H(\Phi(X))$ in IRM leads to $\Phi$ discarding spurious features and grasping invariant features.
Our results empirically support the effectiveness of applying the information bottleneck method to invariant feature learning.
Figure 5: Impact of the information bottleneck on calibration with IB-IRM on CMNIST task. The top row displays the relationship between calibration metrics and Accuracy in the OOD context, investigated by varying the coefficient $\gamma$ of the information bottleneck penalty in the IB-IRM formulation (eq. 9). As the value of $\gamma$ increases, both metrics show improvement. The bottom row visualizes the values of calibration metrics in both ID and OOD. Altering the value of $\gamma$ shows that the larger the value, the more the data aligns with the red line, which indicates equality between the two calibration performances. With a sufficiently large $\gamma$, the points are almost perfectly distributed along the red line, indicating successful calibration across multiple environments.
@@
[linecolor=white, outerlinewidth=1pt, roundcorner=5pt] ID = OOD
@@ 7 Discussion and Conclusion
In this paper, we focus on IRM, which is challenging to implement due to the difficulties of bi-level optimization. We conducted comparative evaluations using ECE, ACE, and NLL, computationally feasible metrics for calibration, to understand its approximation methods better and measure how much the model learned environmental invariant features.
First, our empirical results suggest that calibration performance is equivalent across environments only when calibration performance is relatively high. This suggests that calibration across environments is not only a sufficient condition for invariance, but also a necessary condition empirically, making it clear that calibration is extremely effective for learning invariant features.
It is also suggested that IB-IRM, which applies the information bottleneck method to IRM, achieves more consistent and higher calibration under dataset shift, indicating that it performs invariant feature learning better than other variants.
Another finding from our empirical results is that regularization penalties introduced for OOD generalization effectively improve calibration. However, in more practical datasets (under diversity shift), we observe a trade-off between OOD calibration performance and accuracy. Although IB-IRM mitigates this tradeoff compared to IRMv1 and other variants; carefully balancing these two metrics is important in aiming for OOD generalization.
@@ A.1.1 IRMv1
The learning process was conducted using the loss calculated based on the formulation shown in 8. The hyperparameter $\lambda$ is used to adjust the impact of the IRMv1 penalty.
@@ A.1.2 IBIRM
The learning process was performed using the loss calculated based on the formulation shown in 9. The hyperparameters $\lambda$ and $\gamma$ are used to adjust the influence of the IRMv1 penalty and the information bottleneck penalty, respectively.
@@ Appendix B Additional results
B.1 Ablation Study on Diversity Shift
First, Table 4, similar to Table 2, compares the calibration performance of each method under the same early stopping settings. While Table 2 compares the variance in calibration performance across environments, Table 4 provides a straightforward comparison of the calibration metrics. Although the methods achieve similar OOD accuracy, their calibration performance varies, with IB-IRM achieving consistently better average calibration performance.
Next, Figure 4(c) visualizes the relationship between ID and OOD calibration performance for each dataset. The red solid line represents cases where calibration performance is equal in both domains, indicating successful cross-environment calibration. Notably, IB-IRM (light blue) is distributed relatively close to the red solid line in RMNIST and VLCS, while PAIR (yellow) is close to the red line in PACS. This suggests that these methods achieve consistent calibration across environments.
Combined with the results in Table 2, it can be concluded that IB-IRM consistently achieves high calibration performance across environments on average.
Additionally, from the plots in Figure 4(c), a trend can be observed where lower calibration performance tends to be closer to the red solid line. This implies that achieving consistent calibration across environments, in other words, achieving invariance, requires higher calibration performance.
ERM IRMv1 IB-IRM BIRM IRM Game PAIR
OOD ACC RMNIST 88.3
$\pm$0.1
89.7
$\pm$2.3
85.4
$\pm$1.3
87.6
$\pm$1.1
87.8
$\pm$1.4
82.7
$\pm$1.4
PACS 73.8
$\pm$3.9
76.9
$\pm$1.9
75.5
$\pm$0.8
67.2
$\pm$1.7
72.7
$\pm$1.8
76.0
$\pm$1.6
VLCS 70.0
$\pm$0.8
67.7
$\pm$2.0
69.5
$\pm$0.5
70.2
$\pm$2.0
69.5
$\pm$0.7
67.9
$\pm$0.2
Avg. 77.4
$\pm$1.6
78.1
$\pm$2.1
76.8
$\pm$0.9
75.0
$\pm$1.6
76.7
$\pm$1.3
75.5
$\pm$1.1
OOD ECE RMNIST 5.01
$\pm$1.07
5.74
$\pm$1.42
2.40 $\pm$0.64 4.45
$\pm$0.42
6.71
$\pm$1.09
9.55
$\pm$0.80
PACS 13.05
$\pm$5.21
12.97
$\pm$2.15
11.98
$\pm$0.43
18.61
$\pm$1.80
12.16
$\pm$1.52
7.03 $\pm$3.32
VLCS 12.55
$\pm$1.39
22.79
$\pm$0.85
7.97 $\pm$2.29 9.94
$\pm$2.41
8.78
$\pm$1.23
17.79
$\pm$2.96
Avg. 10.20
$\pm$2.56
13.83
$\pm$1.47
7.45 $\pm$1.12 11.00
$\pm$1.54
9.22
$\pm$1.28
11.45
$\pm$2.36
OOD ACE RMNIST 4.94
$\pm$1.06
5.65
$\pm$1.43
2.26 $\pm$0.59 4.40
$\pm$0.42
6.56
$\pm$1.20
9.49
$\pm$0.82
PACS 12.99
$\pm$5.22
12.91
$\pm$2.09
11.86
$\pm$0.41
18.53
$\pm$1.84
12.06
$\pm$1.61
6.74 $\pm$3.33
VLCS 12.48
$\pm$1.37
22.71
$\pm$0.82
8.07 $\pm$2.16 9.88
$\pm$2.37
8.72
$\pm$1.18
17.79
$\pm$2.96
Avg. 10.14
$\pm$2.55
13.76
$\pm$1.45
7.40 $\pm$1.05 10.94
$\pm$1.54
9.11
$\pm$1.33
11.34
$\pm$2.37
OOD NLL RMNIST 0.41 $\pm$0.03 0.43
$\pm$0.10
0.48
$\pm$0.03
0.42
$\pm$0.04
0.54
$\pm$0.11
0.66
$\pm$0.04
PACS 1.07
$\pm$0.36
1.00
$\pm$0.19
0.87
$\pm$0.03
1.34
$\pm$0.11
1.06
$\pm$0.07
0.75 $\pm$0.05
VLCS 0.97
$\pm$0.08
1.63
$\pm$0.03
0.88 $\pm$0.04 0.90
$\pm$0.08
0.89
$\pm$0.04
1.25
$\pm$0.11
Avg. 0.82
$\pm$0.16
1.02
$\pm$0.11
0.74 $\pm$0.03 0.89
$\pm$0.08
0.83
$\pm$0.07
0.89
$\pm$0.07
Table 4: Comparison of OOD accuracy and the calibration metrics across RMNIST, PACS, and VLCS datasets. Following hyperparameter tuning for each method, measurements were taken across three different random seeds, presenting the mean and variance for each metric. It was observed that on average, IB-IRM exhibits a lower the calibration metrics while maintaining comparable inference performance.
(a) RMNIST
(b) PACS
(c) VLCS
(d) RMNIST
(e) PACS
(f) VLCS
Figure 6: Comparison of the relationship between the calibration metrics in the training environment (horizontal axis) and the test environment (vertical axis). The red solid line represents the case where the calibration metrics is equal in both environments. It was observed that IB-IRM (in light blue) is distributed near the red solid line, indicating a tendency not to overfit to the training environment compared to other methods.
{mdframed}
[linecolor=white, outerlinewidth=1pt, roundcorner=5pt] ERM IRMv1 BIRM IB-IRM IRM Game PAIR ID = OOD
B.2 Impact of penalties of IBIRM
Figure 3 in CMNIST compares the OOD ECE of the model when varying the regularization penalty $\lambda$ in IRMv1. Additionally, we compare the OOD ECE of the model when varying the coefficients of the two regularization penalties in IBIRM Equation 9 as shown in Figure 7 and Figure 8. Similar to Figure 3, as the regularization penalty increases, the OOD ECE decreases.
@@ B.2 Impact of penalties of IBIRM
Figure 7: Comparison of the model’s OOD calibration performance in CMNIST, with various values of $\gamma$ and $\lambda$ fixed at a sufficiently large value (10000) in the formulation of IBIRM 9. It was observed that as $\gamma$ increases, i.e., as the regularization penalty from the information bottleneck method becomes more significant, the calibration performance of the model in OOD scenarios improves.
@@
Figure 8: Comparison of the model’s OOD calibration performance in CMNIST, with various values of $\lambda$ and $\gamma$ fixed at a sufficiently large value (10000) in the formulation of IBIRM 9. It was observed that as $\lambda$ increases, i.e., as the regularization penalty from IRMv1 in IB-IRM becomes more significant, the calibration performance of the model in OOD scenarios improves.
B.3 Feature visualization study
To investigate how differences in ECE impact the model’s ability to learn invariant features, we conducted a feature visualization study. Using a toy CNN, we visualized the latent features of IRMv1 and IB-IRM on RMNIST. The visualizations were created by applying t-SNE (Van der Maaten & Hinton, 2008) to the outputs of the final convolutional layer of the CNN. The results are shown in Figure 9.
First, the table presents the variance in ECE across the six environments of RMNIST, indicating that IB-IRM achieves more stable calibration. Next, consider the four figures. The top row visualizes the feature distributions of each model in the ID setting, while the bottom row shows the OOD setting. The color of each point represents the class, as RMNIST involves 10-class classification.
In a truly successful invariant feature learning scenario, the distributions should be similar in both the top and bottom rows. IB-IRM (right column) shows stable class feature distributions in both ID and OOD. In contrast, IRMv1, with its higher ECE variance, displays variations in these distributions, indicating that the features it extracts differ depending on the environment. This study demonstrates that stable calibration performance across environments reflects true invariance.
B.4 Relationship between ID and OOD accuracy in diversity shift
@@
(a) ID IRMv1(ACC:98.1 ECE:0.008)
(b) ID IB-IRM(ACC:98.1 ECE:0.01)
(c) OOD IRMv1(ACC:92.5 ECE:0.02)
(d) OOD IB-IRM(ACC:92.6 ECE:0.02)
IRMv1 IB-IRM
ECE Variance 2.1e-5 1.0e-5
Figure 9: Visualization of the latent feature spaces for IRMv1 and IB-IRM on RMNIST. The top row shows ID visualizations, and the bottom row shows OOD visualizations. The color of each point represents the class (RMNIST involves 10-class classification). When invariant features are properly acquired, the latent features should remain consistent across environments. IB-IRM, which achieves more stable ECE variance across environments, demonstrates stable feature distributions in both ID and OOD. In contrast, IRMv1, which exhibits worse ECE variance, shows variations in feature distributions between ID and OOD, suggesting it may be acquiring different features depending on the environment.
@@ Appendix C Additional discussion and potential future directions
The experimental results demonstrated that IBIRM, which employs an information bottleneck, excels in terms of model calibration. This can be attributed to the fact that applying the information bottleneck restricts the representational power of the featurizer $\Phi(X)$, limiting its complexity and enabling it to discard spurious features, thereby reducing dependence on them. This approach of constraining the complexity of data representations has been applied in various contexts, and there are related studies in the field of Large Language Models (LLMs) as well as follows.